# Benchmarking Open-ended Audio Dialogue Understanding for Large Audio-Language Models

### v1

@@ 5.3 Results on Each Dataset
Results on ADU-General dataset. The ADU-General dataset aims to evaluate the proficiency in general dialogue understanding, with results across 3 scenarios shown in Fig. 2(a). Our analysis reveals that LALMs perform better in providing helpful responses to helpful questions compared to daily questions and daily statements. Helpful questions typically seek specific information, whereas daily questions and daily statements represent everyday communication between humans, often characterized by a lack of rich contextual information. This finding suggests that LALMs are more adept at handling audio dialogues that require precise information retrieval, while their performance in everyday dialogues remains an area for improvement. In summary, existing open-sourced LALMs understand helpful questions better than daily questions and statements, highlighting the development to better address everyday human interactions.
@@
Results on ADU-Ambiguity dataset. The ADU-Ambiguity dataset is designed to assess how well LALMs handle 4 types of ambiguity during audio dialogue, including intonation-based, pause-based, homophone-based, and repetition-based ambiguity, with results in Fig. 2(d). Overall, LALMs exhibit relatively better performance in handling repetition-based ambiguity, while their performance in managing other types of ambiguities is weaker. This observation suggests that LALMs can more effectively resolve ambiguities that do not involve phonetic elements, such as repetition-based ambiguity, which only has multiple words in an audio. However, when it comes to the other three types of ambiguities, including intonation-based, pause-based, and homophone-based, LALMs struggle to handle them effectively. For homophone-based ambiguity, it is difficult for LALMs to distinguish the words that have almost the same pronunciation. For the other two types of ambiguity, LALMs can not perceive the variations in intonations or pause positions, which can lead to expressing different intentions beyond the same literal meaning of sentences. When faced with these ambiguities, relatively advanced LALMs like GPT-4 often generate responses that encompass multiple possible interpretations. This is due to their inability to effectively distinguish between the different meanings based on phonetic elements. In summary, existing LALMs, including GPT-4 with a Whisper model, display limitations in handling the audio dialogue ambiguity in different phonetic elements, including intonations, pause positions, and homophones.
@@ 5.4 Ablation Study
Human evaluation study. For evaluation, we choose to adopt GPT-4 as the evaluator. To evaluate the consistency between the evaluations of GPT-4 and human judgments, we conduct a human evaluation study as follows. Given the challenge of human testers directly assigning a score on a scale of 0 to 10, we adopt a pairwise comparison approach for models, following (Touvron et al., 2023). Specifically, human testers first listen to the audio queries, then compare two textual responses from two models, finally indicate their preference as “A is better”, “B is better”, or “Both are equal”. We then convert the GPT-4 scores into the same preference-based rating as the human testers. Finally, we evaluate the consistency between the two sets of results, as shown in Fig. 3(b). Our analysis reveals that the pairwise preference consistency achieves a score above 85%, indicating that GPT-4 evaluation aligns well with human judgments. The details are in Appendix D. We provide the evaluation results by LLaMA-3-70B-Instruct and Qwen-2-72B-Instruct and the corresponding human evaluation study in Appendix E.
@@ 6 Conclusion
We present ADU-Bench designed to evaluate the audio dialogue understanding of LALMs. It provides over 20,000 open-ended audio dialogues for LALM assessment for 3 general scenarios, 12 skills, 9 languages, and 4 ambiguity types. Our extensive experiments on 13 LALMs reveal that there is still significant room for improvement in their audio dialogue understanding, which underscores the direction of continued research in LALMs.
@@ Appendix A Generation Details for Synthetic Datasets
[…]
@@
GPT-4 vs BLSP GPT-4 vs SALMONN GPT-4 vs SpeechGPT BLSP vs SALMONN BLSP vs SpeechGPT SALMONN vs PandaGPT
@@ Appendix B Prompts for Evaluation
The evaluation pipeline is shown in Fig. 4. We choose GPT-4 as the default evaluation LLM. We also include LLaMA-3-70B-Instruct and Qwen-2-72B-Instruct to provide the evaluation score. The results are shown in Appendix E.
@@ Appendix C Details of Experimental Settings
To benchmark the audio dialogue understanding of existing LALMs, we assess the performance of 13 LALMs across all 4 datasets within ADU-Bench. Unless stated otherwise, the hyperparameters and setups used during the evaluation process remain consistent with those specified in the original papers of the respective models. For the evaluation, LLaMA-2-7B-Chat, LLaMA-3-8B-Instruct, and LLaMA-3-70B-Instruct are run on 8 NVIDIA A100 40GB GPUs with vLLM library (Kwon et al., 2023), while other open-sourced models are run on a single NVIDIA A100 40GB GPU. By default, our evaluation method employs gpt-4-0613 as the GPT-4 evaluator by calling the API.
@@ Appendix D Human Evaluation Study Details
We conduct a human evaluation study to evaluate the consistency between the evaluations of GPT-4 and human judgments. We also provide a detailed result in Table 3. For the evaluation datasets, we randomly choose 5 audio queries from each domain in ADU-Bench, and finally obtain 140 audio queries. Since ADU-Multilingual contains multiple languages, it is difficult for human testers to understand each language. Hence, we provide the textual transcriptions and allow them to use the translation tools for evaluation. We show them for 10 English speakers. Furthermore, we carefully consider the ethical aspects and potential risks associated with the research involving human subjects. The information we collect is only the preference results and does not involve any personal information.
@@
The instructions given to participants are as follows:
@@
Appendix E Evaluation Results by LLaMA-3-70B-Instruct and Qwen-2-72B-Instruct
@@
Table 4: The average evaluation scores under 13 different LALMs on 4 datasets in our ADU-Bench. The evaluation is conducted by LLaMA-3-70B-Instruct.
@@
Table 5: The average evaluation scores under 13 different LALMs on 4 datasets in our ADU-Bench. The evaluation is conducted by Qwen-2-72B-Instruct.
@@
GPT-4 vs BLSP GPT-4 vs SALMONN GPT-4 vs SpeechGPT BLSP vs SALMONN BLSP vs SpeechGPT SALMONN vs PandaGPT
@@
GPT-4 vs BLSP GPT-4 vs SALMONN GPT-4 vs SpeechGPT BLSP vs SALMONN BLSP vs SpeechGPT SALMONN vs PandaGPT
@@
Appendix F Limitations
The primary limitation of this work lies in the analysis being restricted to a limited number of LALMs, due to constraints such as the availability of usable code, model weights, and the resources required for extensive experimentation. Future work could address this by exploring a broader and more diverse range of LALMs to deepen the analysis. Moreover, ADU-Bench currently focuses on evaluating performance in areas such as daily dialogues, skill capabilities, multilingual proficiency, and natural robustness in handling ambiguities. In the future, we aim to expand ADU-Bench to include additional evaluation domains. For example, if LALMs are integrated into embodied AI systems, it will be crucial to emphasize security-related evaluations. Concretely, we leave the investigation whether LALMs are vulnerable to adversarial misdirection when processing partially corrupted audio inputs (Gu and Tresp, 2020; Gu et al., 2022; Gu, 2024; Li et al., 2024; Gao et al., 2024a) in the future work. Additionally, we will explore how LALMs respond to backdoor attacks, such as when audio clips containing hidden triggers are injected into poisoned datasets (Gao et al., 2023b, a). Another important aspect will be evaluating the models’ behavior under maliciously manipulated energy and latency conditions (Gao et al., 2024b, c, d), examining whether they can be forced into infinite loops that unnecessarily consume resources. By incorporating a wider variety of LALMs and extending the coverage of dialogue domains, we aim to ensure that ADU-Bench remains comprehensive and up-to-date.
Appendix G Broader Impacts
@@ Appendix I More Overall Results
Table 9: The score for audio dialogue understanding performances under 13 different LALMs on 4 datasets in our proposed ADU-Bench. The textual reference is before the textual response in the evaluation prompt for the GPT-4 evaluator.
@@
Table 10: The score for audio dialogue understanding performances under 13 different LALMs on 4 datasets in our proposed ADU-Bench. The textual reference is after the textual response in the evaluation prompt for the GPT-4 evaluator.
@@
Table 11: The score for audio dialogue understanding performances under 13 different LALMs on ADU-General dataset. The textual reference is before the textual response in the evaluation prompt for the GPT-4 evaluator.
@@
Table 12: The score for audio dialogue understanding performances under 13 different LALMs on ADU-General dataset. The textual reference is after the textual response in the evaluation prompt for the GPT-4 evaluator.
@@
Table 13: The score for audio dialogue understanding performances under 13 different LALMs on ADU-Skill dataset. The textual reference is before the textual response in the evaluation prompt for the GPT-4 evaluator.
@@
Table 14: The score for audio dialogue understanding performances under 13 different LALMs on ADU-Skill dataset. The textual reference is before the textual response in the evaluation prompt for the GPT-4 evaluator.
@@
Table 15: The score for audio dialogue understanding performances under 13 different LALMs on ADU-Skill dataset. The textual reference is after the textual response in the evaluation prompt for the GPT-4 evaluator.
@@
Table 16: The score for audio dialogue understanding performances under 13 different LALMs on ADU-Skill dataset. The textual reference is after the textual response in the evaluation prompt for the GPT-4 evaluator.
@@
Table 17: The score for audio dialogue understanding performances under 13 different LALMs on ADU-Multilingual dataset. The textual reference is before the textual response in the evaluation prompt for the GPT-4 evaluator.
@@
Table 18: The score for audio dialogue understanding performances under 13 different LALMs on ADU-Multilingual dataset. The textual reference is before the textual response in the evaluation prompt for the GPT-4 evaluator.
@@
Table 19: The score for audio dialogue understanding performances under 13 different LALMs on ADU-Multilingual dataset. The textual reference is after the textual response in the evaluation prompt for the GPT-4 evaluator.
@@
Table 20: The score for audio dialogue understanding performances under 13 different LALMs on ADU-Multilingual dataset. The textual reference is after the textual response in the evaluation prompt for the GPT-4 evaluator.
@@
Table 21: The score for audio dialogue understanding performances under 13 different LALMs on ADU-Ambiguity dataset. The textual reference is before the textual response in the evaluation prompt for the GPT-4 evaluator.
@@
Table 22: The score for audio dialogue understanding performances under 13 different LALMs on ADU-Ambiguity dataset. The textual reference is after the textual response in the evaluation prompt for the GPT-4 evaluator.

### v2

@@ 5.3 Results on Each Dataset
Results on ADU-General dataset. The ADU-General dataset aims to evaluate the proficiency in general dialogue understanding, with results across 3 scenarios shown in Fig. 2(a). Our analysis reveals that LALMs perform better in helpful questions compared to daily questions and daily statements. Helpful questions typically seek specific information, whereas daily questions and daily statements represent everyday communication between humans, often characterized by a lack of rich contextual information. This finding suggests that LALMs are more adept at handling audio dialogues that require precise information retrieval, while their performance in everyday dialogues remains an area for improvement. In summary, existing open-sourced LALMs understand helpful questions better than daily questions and statements, highlighting the continued development in LALMs to better address everyday human interactions.
@@
Results on ADU-Ambiguity dataset. The ADU-Ambiguity dataset is designed to assess how well LALMs handle 4 types of ambiguity during audio dialogue, including intonation-based, pause-based, homophone-based, and repetition-based ambiguity, with results in Fig. 2(d). Overall, LALMs exhibit relatively better performance in handling repetition-based ambiguity, while their performance in managing other types of ambiguities is weaker. This observation suggests that LALMs can more effectively resolve ambiguities that do not involve phonetic elements, such as repetition-based ambiguity, which only has multiple words in an audio. However, when it comes to the other three types of ambiguities, including intonation-based, pause-based, and homophone-based, LALMs struggle to handle them effectively. For homophone-based ambiguity, it is difficult for LALMs to distinguish the words that have almost the same pronunciation. For the other two types of ambiguity, LALMs can not perceive the variations in intonations or pause positions, which can lead to expressing different intentions beyond the same literal meaning of sentences. When faced with these ambiguities, relatively advanced LALMs like GPT-4o can achieve an average score of 5.22 and 6.05 for pause-based and homophone-based ambiguity. The results show that GPT-4o often generates responses that encompass multiple possible interpretations, which is unable to effectively distinguish between the different meanings based on phonetic elements. In summary, existing LALMs, including GPT-4o, display limitations in handling the audio dialogue ambiguity in different phonetic elements.
@@ 5.4 Ablation Study
Human evaluation study. For evaluation, we choose to adopt GPT-4 as the evaluator. To evaluate the consistency between the evaluations of GPT-4 and human judgments, we conduct a human evaluation study as follows. Given the challenge of human testers directly assigning a score on a scale of 0 to 10, we adopt a pairwise comparison approach for models, following (Touvron et al., 2023). Specifically, human testers first listen to the audio queries, then compare two textual responses from two models, finally indicate their preference as “A is better”, “B is better”, or “Both are equal”. We then convert the GPT-4 scores into the same preference-based rating as the human testers. Finally, we evaluate the consistency between the two sets of results, as shown in Fig. 3(b). Our analysis reveals that the pairwise preference consistency achieves a score above 85%, indicating that GPT-4 evaluation aligns well with human judgments. The details are in Appendix D. We provide the evaluation results by LLaMA-3-70B-Instruct and Qwen-2-72B-Instruct and the corresponding human evaluation study in Appendix F.
@@ 6 Conclusion
In conclusion, we present ADU-Bench, a comprehensive benchmark designed to evaluate the audio dialogue understanding of LALMs. It encompasses 4 benchmark datasets including ADU-General dataset for 3 general scenarios, ADU-Skill dataset for 12 skills, ADU-Multilingual dataset for 9 languages, and ADU-Ambiguity for 4 ambiguity types, providing over 20,000 open-ended audio dialogues for the LALM assessment. Our extensive experiments on 16 LALMs reveal that there is still significant room for improvement in their audio dialogue understanding. Notably, LALMs face challenges in processing mathematical symbols and formulas, comprehending human behavior like roleplay, understanding multiple languages, and handling audio dialogue ambiguities arising from various phonetic elements.
Limitations
The main limitation of this work is that the analysis is on a limited number of LALMs due to the availability of usable code, model weights, and massive experiments. Potential future work includes investigating more diverse LALMs and designing more domains about audio dialogues to make our ADU-Bench up-to-date.
Ethics Statement
Acknowledgement
This work is supported in part by the National Natural Science Foundation of China under Grant 62171248 and Shenzhen Science and Technology Program (JCYJ20220818101012025).
@@ Appendix A Generation Details for Synthetic Datasets
To generate synthetic datasets for ADU-General dataset, ADU-Skill dataset, and ADU-Multilingual dataset, we first adopt GPT-4 and human inspection to obtain the related textual dialogues for each dataset. Then, enclose them in the Speech Synthesis Markup Language (SSML) (Taylor and Isard, 1997) by human coding, where SSML is an XML-based markup language specifically designed for speech synthesis applications. Subsequently, execute the program code using Python interpreter with public SSML service (Microsoft, 2024) provided by Microsoft Azure to convert them into audio dialogues. Furthermore, to emulate real-world scenarios, we consider a wide array of variables for synthetic audio. They include 2 genders (male and female), 4 different speakers (2 men and 2 women), 4 emotions (calm, excited, angry, and sad), 3 speech rates (standard and $\pm 10\%$± 10 %), 3 pitch levels (standard and $\pm 10\%$± 10 %), and 3 volume levels (standard and $\pm 10\%$± 10 %). During the generation of each dataset, a combination of these audio generation characteristics is randomly selected to create each audio data, ensuring diversity in the audio dialogues. Therefore, this generation method not only provides a scalable solution for generating synthetic audio datasets but also ensures a rich diversity that closely mirrors real-world audio dialogues.
@@
GPT-4 vs
BLSP
GPT-4 vs
SALMONN
GPT-4 vs
SpeechGPT
BLSP vs
SALMONN
BLSP vs
SpeechGPT
SALMONN vs
PandaGPT
@@ Appendix B Prompts for Evaluation
The evaluation pipeline is shown in Fig. 4. We choose GPT-4 as the default evaluation LLM. We also include LLaMA-3-70B-Instruct and Qwen-2-72B-Instruct to provide the evaluation score. The results are shown in Appendix F.
@@ Appendix C Details of Experimental Settings
To benchmark the audio dialogue understanding of existing LALMs, we assess the performance of 16 LALMs across all 4 datasets within ADU-Bench. Unless stated otherwise, the hyperparameters and setups used during the evaluation process remain consistent with those specified in the original papers of the respective models. For the evaluation, LLaMA-2-7B-Chat, LLaMA-3-8B-Instruct, and LLaMA-3-70B-Instruct are run on 8 NVIDIA A100 40GB GPUs with vLLM library (Kwon et al., 2023), while other open-sourced models are run on a single NVIDIA A100 40GB GPU. By default, our evaluation method employs gpt-4-0613 as the GPT-4 evaluator by calling the API.
@@ Appendix D Human Evaluation Study Details
We conduct a direct human evaluation on randomly selected 140 audio dialogues from ADU-Bench. Each sample is assessed by three human testers, who rate the generated responses. Human testers should provide an overall score on a scale of 0 to 10, where a higher score indicates better overall performance. The results are shown in Table 2. Besides, we conduct a human evaluation study to evaluate the consistency between the evaluations of GPT-4 and human judgments. We show each pair of samples for ten human testers. The results are demonstrated in Fig. 3(b) and Table 3.
For the evaluation datasets, we randomly choose 5 audio queries from each domain in ADU-Bench, and finally obtain 140 audio queries. Since ADU-Multilingual contains multiple languages, it is difficult for human testers to understand each language. Hence, we provide the textual transcriptions and allow them to use the translation tools for evaluation. we carefully consider the ethical aspects and potential risks associated with the research involving human subjects. The information we collect is only the preference results and does not involve any personal information.
@@
The instructions given to participants in Table 2 are as follows:
In this study, you will be presented with a total of 140 audio clips, each accompanied by one textual response. For audio in foreign languages, we will provide textual transcriptions and translation tools to assist you.
Your task is as follows:
1. Listen to the audio queries carefully.
2. Based on the criteria of helpfulness, relevance, accuracy, and comprehensiveness, provide an overall score on a scale of 0 to 10 for the response, where a higher score indicates better overall performance.
The instructions given to participants in Fig. 3(b) and Table 3 are as follows:
@@
Appendix E Discussions
E.1 Real-world and Synthetic Audio in ADU-Bench
Our ADU-Bench includes both real-world and synthetic audio. As stated in Section 3, the data collection involves a combination of synthetically generated dialogues and real-world audio samples. Specifically, our ADU-Bench contains over 8000 audio samples from the real world. A reason prevents us from using real-world audio only is the challenges and costs of the collection process. In particular, the collection of professional technical terms or languages can be difficult, as it requires humans who are familiar with them. Without proper familiarity, the use of these terms or languages in audio samples may sound unnatural. To address this issue, we propose a synthetic method for audio generation in Appendix A. By leveraging it, we can easily expand a scalable ADU-Bench without incurring substantial expenses. Besides, we have conducted an ablation study to investigate the effects of real-world and synthetic audio on the performance of our benchmark, as detailed in Section 5.4. It can be observed that there is no significant difference in the performance of LALMs in the areas our ADU-Bench covers. It illustrates that these synthetic audios can also benchmark audio dialogue understanding.
E.2 Evaluation using Textual Response
In our evaluation process, we prompt audio queries to obtain audio responses and adopt their textual transcriptions with references to calculate a GPT-4 evaluation score. We have chosen this approach because our primary focus is on how LALMs comprehend audio dialogue and formulate appropriate replies. In our ADU-Bench, we emphasize understanding ability rather than generation quality of audio dialogues. Directly using audio for evaluation can be challenging, and evaluating generation quality is not within the scope of ADU-Bench. By opting for a textual format, we can concentrate on assessing LALMs’ dialogue understanding abilities and their capacity to provide meaningful responses, without introducing the additional complexity of audio generation. Furthermore, our evaluation approach in ADU-Bench aligns with previous work (Yang et al., 2024).
E.3 Analysis for Weak Performance of LALMs
LALMs consist of two main components - audio feature extractors and base LLMs. For textual benchmarks such as GSM8K and MMLU, the base LLMs of LALMs are usually able to achieve effective mathematical and knowledge-based reasoning, which reflects the fundamental reasoning skills of the base LLMs. However, for our ADU-Bench, the LALMs exhibit weak performance and are unable to demonstrate the fundamental reasoning skills of their base LLM components. This observation leads us to conjecture that the poor performance of the LALMs on the ADU-Bench is primarily rooted in their audio comprehension abilities, rather than their core reasoning skills.
Appendix F Evaluation Results by LLaMA-3-70B-Instruct and Qwen-2-72B-Instruct
@@
Table 4: The average evaluation scores under 16 different LALMs on 4 datasets in our ADU-Bench. The evaluation is conducted by LLaMA-3-70B-Instruct.
@@
GPT-4o - 8.69 8.35 8.61 6.37 8.00
Table 5: The average evaluation scores under 16 different LALMs on 4 datasets in our ADU-Bench. The evaluation is conducted by Qwen-2-72B-Instruct.
@@
GPT-4 vs
BLSP
GPT-4 vs
SALMONN
GPT-4 vs
SpeechGPT
BLSP vs
SALMONN
BLSP vs
SpeechGPT
SALMONN vs
PandaGPT
@@
GPT-4 vs
BLSP
GPT-4 vs
SALMONN
GPT-4 vs
SpeechGPT
BLSP vs
SALMONN
BLSP vs
SpeechGPT
SALMONN vs
PandaGPT
@@
Appendix G Reproducibility Statement
We provide the code and data in the project page of our ADU-Bench.
@@ Appendix I More Overall Results
Table 9: The score for audio dialogue understanding performances under 16 different LALMs on 4 datasets in our proposed ADU-Bench. The textual reference is before the textual response in the evaluation prompt for the GPT-4 evaluator.
@@
GPT-4o - 8.54 8.84 8.07 6.79 8.06
Table 10: The score for audio dialogue understanding performances under 16 different LALMs on 4 datasets in our proposed ADU-Bench. The textual reference is after the textual response in the evaluation prompt for the GPT-4 evaluator.
@@
Table 11: The score for audio dialogue understanding performances under 16 different LALMs on ADU-General dataset. The textual reference is before the textual response in the evaluation prompt for the GPT-4 evaluator.
@@
   GPT-4o    -    8.76    8.65    8.20
Table 12: The score for audio dialogue understanding performances under 16 different LALMs on ADU-General dataset. The textual reference is after the textual response in the evaluation prompt for the GPT-4 evaluator.
@@
   GPT-4o    -    8.92    8.74    8.55
Table 13: The score for audio dialogue understanding performances under 16 different LALMs on ADU-Skill dataset. The textual reference is before the textual response in the evaluation prompt for the GPT-4 evaluator.
@@
GPT-4o - 9.53 9.06 8.67 8.98 9.21 8.84
Table 14: The score for audio dialogue understanding performances under 16 different LALMs on ADU-Skill dataset. The textual reference is before the textual response in the evaluation prompt for the GPT-4 evaluator.
@@
GPT-4o - 8.41 9.25 8.32 8.71 8.14 8.98
Table 15: The score for audio dialogue understanding performances under 16 different LALMs on ADU-Skill dataset. The textual reference is after the textual response in the evaluation prompt for the GPT-4 evaluator.
@@
GPT-4o - 9.76 9.35 8.84 9.12 9.36 9.03
Table 16: The score for audio dialogue understanding performances under 16 different LALMs on ADU-Skill dataset. The textual reference is after the textual response in the evaluation prompt for the GPT-4 evaluator.
@@
GPT-4o - 8.74 9.38 8.54 9.14 8.87 9.12
Table 17: The score for audio dialogue understanding performances under 16 different LALMs on ADU-Multilingual dataset. The textual reference is before the textual response in the evaluation prompt for the GPT-4 evaluator.
@@
   GPT-4o    -    7.40    6.80    9.09    9.25    8.69
Table 18: The score for audio dialogue understanding performances under 16 different LALMs on ADU-Multilingual dataset. The textual reference is before the textual response in the evaluation prompt for the GPT-4 evaluator.
@@
  GPT-4o   -   7.34   7.28   8.01   8.73
Table 19: The score for audio dialogue understanding performances under 16 different LALMs on ADU-Multilingual dataset. The textual reference is after the textual response in the evaluation prompt for the GPT-4 evaluator.
@@
   GPT-4o    -    7.51    7.12    9.24    9.35    8.84
Table 20: The score for audio dialogue understanding performances under 16 different LALMs on ADU-Multilingual dataset. The textual reference is after the textual response in the evaluation prompt for the GPT-4 evaluator.
@@
  GPT-4o   -   7.58   7.43   8.21   8.85
Table 21: The score for audio dialogue understanding performances under 16 different LALMs on ADU-Ambiguity dataset. The textual reference is before the textual response in the evaluation prompt for the GPT-4 evaluator.
@@
GPT-4o - 6.96 5.11 5.97 9.10
Table 22: The score for audio dialogue understanding performances under 16 different LALMs on ADU-Ambiguity dataset. The textual reference is after the textual response in the evaluation prompt for the GPT-4 evaluator.